A ball is thrown vertically upwards with a velocity of 10 m/s.It returns to the ground with a velocity of 9 m/s.What is the height attained by the ball ?

Grade:9

3 Answers

SURESH
8 years ago
dear vijaya patel,
                          we know the equation of height h=u^2/2g (g=gravity of earth is 9.8 m/s^2) & u is initial velocity i.e.,10 m/s we caliculate by this formula we get 5.102 m so, the ball reached 5.102 m.
Vijeya Patel
8 years ago
But in both the cases the distance is different sir.(when going up and coming down).What about that?and my textbook gives th answer 4.62m ,but I don’t understand how?please help
NARAYANA
8 years ago
WE KNOW THE EQUATION FOR HIGHT ANY OBJECT THROWN WITH INTIAL VELOCITY(U),GRAVITY(g)THEN EQUATION IS U^2/2g,SO ANSWER IS 10^2/2*9.8=5.10M